The perimeter of a triangle in which the lengths of all of the sides are integers is 21. If the length of one side of the triangle is 8, what is the shortest possible length of another side of the triangle
Answer by checkley77(12844)

21=A+B+C
21=8+B+C
21-8=B+C
13=B+C
IF THE SHORT SIDE WAS 1 THE OTHER SIDE WOULD HAVE TO BE 12. WHICH WOULD NOT A TRIANGLE MAKE. (8+1<12)
IF THE SHORT SIDE WAS 2 THE OTHER SIDE WOULD HAVE TO BE 11. AGAIN NOT A TRIANGLE. (8+2<11)
IF THE SHORT SIDE WAS 3 THE OTHER SIDE WOULD HAVE TO BE 10. THESE SIDES WOULD MAKE A TRIANGLE. (8+3>10)
ANSWER IS 3 FOR THE SHORTEST SIDE OF THE TRIANGLE.